OpenStudy (anonymous):

How do I solve this equation: sin( (2π/3)-x ) = sin( x ) (0<=x<=7π/3)

OpenStudy (jhannybean):

\[\sf \begin{align*} \sin\left(\frac{2\pi}{3} -x\right)=\sin(x)\\ &:~ \color{red}{\sin(\alpha-\beta) = \sin(\alpha)\cos(\beta)-\cos(\alpha)\sin(\beta)}\\ &: ~ \color{blue}{\sin\left(\frac{2\pi}{3}\right)}\cos(x)-\color{blue}{\cos\left(\frac{2\pi}{3}\right)}\sin(x)=\sin(x)\\ &:\frac{\sqrt{3}}{2}\cos(x)+\left(\frac{1}{2}\right)\sin(x)=\sin(x)\\ &: ~ \frac{\sqrt{3}}{2}\cos(x) = \frac{1}{2}\sin(x)\\ \end{align*}\]

OpenStudy (jhannybean):

Ah... hmm... We want constants on one side, but our function on the other. We can divide both sides by either cosine or sine. I'll divide both sides by cos(x)

OpenStudy (jhannybean):

\[\begin{align*} \frac{\sqrt{3}}{2}\cos(x)=\frac{1}{2}\sin(x) \\&: ~ \sqrt{3}=\tan(x) \\&: ~ \tan^{-1}(\sqrt{3} ) = \tan^{-1}(\tan(x)) \\&: ~ \tan^{-1}(\sqrt{3}) = x \\&: ~ x=\frac{\pi}{3} \end{align*}\]Since we are trying to find all the values between \(0\) and \(\dfrac{7\pi}{3}\) We need to find the general solution for x, For tangent functions, we know the period is \(\pi\) so starting at \(\dfrac{\pi}{3}\) and adding it to our period, \(k\pi\) where \(k=1,2,3...\) We can state that: \(\boxed{x=\dfrac{\pi}{3} +k\pi} \) is our solution.

OpenStudy (cwrw238):

hmm... try expanding LHS: sin (2pi/3) cos x - cos (pi/3) sinx = sin x

OpenStudy (cwrw238):

sin 2pi/3 = sqrt3/2 and cos 2pi/3 = -1/2

OpenStudy (cwrw238):

giving us (qrt3/2) cos x - (1/2) sin x = 0

OpenStudy (cwrw238):

(sqrt3 / 2) cos x = (1/2) sin x divide both sides by cos x sqrt3/2 = 1/2 tan x solve for x

OpenStudy (mathmate):

Another way: From \(\large \sin\left(\frac{2\pi}{3} -x\right)=\sin(x)\\\) we deduce \(\large \left(\frac{2\pi}{3} -x\right)=x+2k\pi\\\) where \(k\in Z\) Solve for x \(\large x=\frac{\pi}{3}+k\pi\) as before, but limit values of k for \([0\le x \le 7\pi/3]\)